What is 3v+7/-8=20
Answer by TimothyLamb(4379)

3v + 7/-8 = 20
3v - 7/8 = 20
3v = 20 + 7/8
3v = 160/8 + 7/8
3v = 167/8
---
v = (1/3)(167/8)
---
Answer:
v = 6.9583333
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php